OpenStudy (anonymous):

what is the solution of the following system x-y=11 -x+y=-11

OpenStudy (anonymous):

0 = 0 infinitely many solutions

hero (hero):

x - y = 11 -x + y = -11 Divide both sides of 2nd equation by -1 x - y = 11 x - y = 11 Both equations are the same. That's why there are infinite solutions.